OpenStudy (anonymous):

A line passes through the point (0, 2) and has a slope of -1/4. What is the equation of the line? A. 4x + y = 2 B. x - 4y = 8 C. 4x - y = -2 D. x + 4y = 8 I know the slope-intercept form would be y = -1/4 + 2, but I don't know how to convert it to standard.

OpenStudy (tkhunny):

That is not the Slope-Intercept form. Seems to be an 'x' missing. y = -(1/4)x + 2 Have you considered multiplying by 4?

OpenStudy (anonymous):

Sorry, I didn't mean to leave it out. Would you multiply each one by 4?

OpenStudy (anonymous):

The standard form of a line is : ax+by=C So basically our first step would be to bring the x and the y to the same side and the C which is in this case 2 to the other side of the equation

OpenStudy (anonymous):

so the easiest way to do this would be to add (1/4)x to both sides y+(1/4)x=(1/4)x-(1/4)x+2 y+(1/4)x=2

OpenStudy (anonymous):

Now we dont want to have fractions in our equation so in order to gid rid of the fraction we mulitply the whole equation by 4 4*y+4*(1/4)x=2*4 4y+x=8

OpenStudy (anonymous):

Thanks so much. I've finally got it

OpenStudy (anonymous):

Great :)